OpenStudy (anonymous):

need a little help f(x)=5-3x for x=1 and x=3 would you plug the x=1 in for the f(x) or the 3x

OpenStudy (anonymous):

replace all the x's with a 1

OpenStudy (anonymous):

well whats the x=3 for

OpenStudy (anonymous):

f(x)=5-3x f(1) = 5-3(1)

OpenStudy (anonymous):

some with the 3

OpenStudy (anonymous):

it's two separate questions

OpenStudy (anonymous):

x can't be both 1 and 3 at the same time

OpenStudy (anonymous):

so does it equal to 3 ?

OpenStudy (anonymous):

f(x)=5-3x is a function with x as a variable. the outcome of the function changes depending on the input. the input is the value for x you choose, the output is the result of the equation

OpenStudy (anonymous):